If the atomic weight of an element is 23 times that of the lightest element and it has 11 protons then its contains

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi pallavi,

atomic weight = 23 * atomic weight of hydrogen = 23*1 =23

given it has 11 protons....implies no of neutrons=23-11=12